OpenStudy (anonymous):

Having received his weekly allowance, John spent 3/5 of his allowance at the arcade. The next day he spent one third of his remaining allowance at the toy store, and then spent his last $0.80 at the candy store. What is John’s weekly allowance?

OpenStudy (anonymous):

if his allowance is x, 3x/5 = allowance on arcade this means that there is 2x/5 remaining 1/3(2x/5) = toy store => 2x/15 = toy store adding 3/5 and 2/15 we get a total amount spent of 11x/15 of his allowances. this must mean that there is only 4x/15 remaining. since his last bit of money was .80, we can say 4x/15 = .80. solving for x we have x = $3.00

OpenStudy (anonymous):

why wouldnt it be 1/3x (2/5x)

OpenStudy (anonymous):

and y do u multiply instead of adding

OpenStudy (anonymous):

because read the question again, he spent 3/5 at the arcade, then he spent 1/3 (emphasis on THIRD) on his REMAINING allowance, not 1/5 of his total allowance, which is why you multiply

OpenStudy (anonymous):

i dont get it can u give me a visual and draw something?

OpenStudy (anonymous):

hey y do u have it multiplied by 1/3 instead of 2/3?

OpenStudy (anonymous):

because the question says he spent 1/3 of his remaining at the toy store not 2/3

OpenStudy (anonymous):

but isnt the remaining 2/3?

OpenStudy (anonymous):

we use what he spent, not what is remaining

OpenStudy (anonymous):

he spent 1/3 of his allowance AFTER he spent the 3/5

OpenStudy (anonymous):

you dont use the remaining allowance to calculate

OpenStudy (anonymous):